# Also at ] The European Space Agency Centre for Earth Observation, Frascati, Italy

### v3

@@ Abstract
Recent minimalist modeling has demonstrated that overdamped polar chiral active matter can support emergent, inviscid Euler turbulence, despite the system’s strictly dissipative microscopic nature. In this article, we establish the statistical mechanical foundation for this emergent inertial regime by deriving a formal isomorphism between the model’s agent dynamics and the overdamped Langevin equation for disordered Josephson junctions. We identify the trapped agent state as carrying non-dissipative phase rigidity supercurrents, a mapping we confirm empirically by demonstrating a disorder-broadened Adler-Ohmic crossover in the system’s slip velocity. Generalizing this framework to three dimensions ($S^{1}\to S^{2}$), we show that the polar alignment interaction is geometrically equivalent to the Gilbert damping term in spintronics. This constraint naturally generates inertial spin waves (ferromagnetic magnons) from the overdamped active bath, recovering the macroscopic transport predicted by Toner-Tu theory without invoking microscopic inertia. These results classify chiral active matter as a dissipative spintronic fluid, revealing that the transition to active turbulence is the emergence of a topological quasiparticle gas, where macroscopic Onsager dipoles acquire effective inertia from the Goldstone modes of the underlying broken symmetry.
@@ I Introduction
A recent minimalist, two-dimensional chiral active matter model offers a solution to the paradox by empirically demonstrating that an overdamped chiral flock can drive a conservative inertial cascade that culminates in inviscid Euler turbulence, in the form of a stable, sloshing Onsager dipole in the dynamic agent ensemble, where topological defects interact long-range through the Kirchoff-Onsager hamiltonian [8] (Figure 1). The present article provides detailed theoretical backdrop for the findings in Ref. [8].
@@
The model is characterized by three simple rules, namely that its agents are (1) self-propelled, (2) polar, or flocking, favoring alignment, and (3) chiral, with intrinsic rotation [12, 8]. The model is unique in that it collapses these three properties into a single degree of freedom, phase, $\phi_{i}$,
@@
However, an important question remains. Our active agent ensemble is demonstrably overdamped and dissipative, yet the model exhibits an inverse, conservative energy cascade. How does a system that should fundamentally decay instead pump energy towards the largest scales? The resolution lies in considering the discrete distribution of agents as a spatially extended, disordered Josephson junction array. The depinning transition between ’running’ and ’trapped’ states then generates a macroscopic synchronization stiffness, or supercurrents of phase gradient information. By extending the model’s equations to three dimensions (3D), we eventually demonstrate that the rigidity supports Goldstone mode spin waves that map the two-dimensional (2D) model to the azimuthal precession of 3D active chiral agents. We shall contrast these ferromagnetic magnons against the second sound waves of the Toner-Tu field theory, delineating the universality of the Landau-Lifshitz-Gilbert equation.
@@ IV Reconciling the model with real, three-dimensional chiral flocks
$\bm{\Omega}_{slip}=\underbrace{(\bm{\omega}_{i}-\bm{\Omega}_{\Psi})}_{\equiv\;\mathbf{I}_{eff}}-\underbrace{a_{0}R(\hat{\mathbf{\Psi}}\times\hat{\mathbf{n}}_{i})}_{\equiv\;\mathbf{I}_{c}}.$ (22)
@@
That is, the minimal agent dynamics mapping a two-dimensional unit phase vector to a three-dimensional unit vector naturally leads to an Adler relation. Remarkably, this occurs without the explicit definition of a Kuramoto-term. The reason for this stem from the recognition of the Landau-Lifshitz-Gilbert-equation from ferromagnetic theory, where a spin $\mathbf{S}$ naturally relaxes towards a magnetic field $\mathbf{H}$ (Figure 4b). The three-dimensional synchronization in a chiral flock described in Eqs. (19) is therefore equivalent to magnetic damping.
@@ V Overdamped spin waves as ‘second sound’ in Toner-Tu theory
The previous section established a three-dimensional adherence to the Adler equations, where flock synchronization naturally arises from a Kuramoto-like torque forced by the geometry of $S^{2}$ (Eq. 19). Having established the rigidity of the flock, the natural next step is to demonstrate that the predicted effects are observable in real 3D flocks. To do so, we must reconcile the 3D extension of our minimalist model with the standard model invoked to explain flocking in three dimensions: the Toner-Tu framework [9]. We begin with the 3D equations of motion, Eq. (19). To study the stability of the ordered state, we simplify by neglecting the intrinsic chirality $\bm{\omega}_{i}$, focusing on strong alignment:
@@
where $\xi$ is the characteristic interaction scale. This approximation follows from the definition of $\hat{\Psi}$ as the average director within an interaction radius $\sigma$. Treating the agent distribution as a continuum, $\hat{\Psi}(\mathbf{r})$ is the convolution of $\hat{n}(\mathbf{r})$ with a local weighting kernel. A Taylor expansion of $\hat{n}(\mathbf{r}+\bm{\delta})$ averaged over an isotropic neighborhood eliminates first-order gradient terms ($\langle\bm{\delta}\rangle=0$). The leading-order correction is therefore the quadratic term, proportional to $\nabla^{2}\hat{n}$, which encodes the local field curvature. Consequently, $\xi^{2}$ scales with the square of the interaction radius ($\xi\sim\sigma^{2}$). Substituting these continuum relations into the alignment term of Eq. (26) yields:
@@
This form is physically significant as it mirrors the dissipative damping term $\hat{n}\times(\hat{n}\times\mathbf{H})$ found in the Landau-Lifshitz-Gilbert equation of spintronics [34]. To fully capture the wave-like transport observed in Toner-Tu theories [9], we restore the inertial term neglected in the overdamped limit. We introduce a coefficient of turning inertia, or precession stiffness, $\gamma$:
@@
We note that this restoration of effective inertia is physically mandated by the conservation of angular momentum density: as the local order parameter $R$ vanishes, the collective moment of inertia disappears ($I_{eff}\propto R$), compelling the precession rate to diverge to sustain a finite interaction torque. This description remains valid only while the local timescale separation persists, breaking down precisely at the defect cores ($R\rightarrow 0$) where the divergence signals the transition from a hydrodynamic superfluid to a strictly microscopic, overdamped gas. To satisfy dimensional consistency ($\gamma]=L^{2}T^{-1}$) within this physical constraint, we define the inertial coefficient as:
@@
which leads to a direct mapping to the sound speed of the topological gas, $c_{s}\approx v_{0}R$,
@@
Physically, the inverse dependency in Eq. (31) implies that at the flock boundary ($R\to 0$), as alluded to, the precession frequency diverges. The system transitions from a “heavy,” locked superfluid in the bulk to a massless, fast-spinning gas at the edges. This divergence creates the active bath required to sustain the topological protection of the inner Onsager dipole [8].
@@
The active, 3D, overdamped bath on the boundaries of a real chiral flock naturally leads to a conservative azimuthal turning inertia of a spintronic fluid [34], by merit of the dimensional constraints of chiral alignment alone. The conservative fluid forms the characteristic flocking behaviour, and is strictly contingent on the agents individually adhering to the phase-locking condition from Josephson junction theory (Eq. 25). The conservation is the relaxation of the stress-energy of the sphere $S^{2}$, giving a geometric and thereby completely natural definition of synchronization in any system with polar alignment, by merit of the Landau-Lifshitz-Gilbert equation equation.
@@ VI.2 Thermodynamic pump
To see why, we identify the azimuthal angle of the $S^{2}$ potential as the scalar phase $\phi$ in the reduced $S^{1}$ limit. This degree of freedom is driven by the intrinsic chirality $\omega_{i}$ [and stochastic noise $\eta_{i}(t)$], providing the essential ”frustration” mandated by Ref. [24] that prevents the chiral flock from freezing into a trivial, fully organized state. Instead, the frustration fuels a continuous, natural precession within the washboard potential wells (Figure 5b), generating the Goldstone mode quasiparticles (magnons) that sustain the fluid’s inertia.
@@
Inside the two-dimensional Onsager dipole’s vortex cores, we observe in Figure 2a,d) that the agents are continuously phase-slipping. These cores are identified by Ref. [8] as massless ($R\to 0$) sonic black holes. In this limit, the critical current Eq. (12) vanishes, reducing the Adler equation to its trivial kinematic limit ($\dot{\phi}_{i}\approx\omega_{i}$), and the spintronic fluid’s capacity to support spin waves diverges, $\gamma\propto R^{-1}$ (Eq. 31). Consequently, the vortex cores are purely dissipative singularities where the capacity to sustain supercurrents has collapsed. This decoupling renders the cores unable to transmit phase stiffness or radiate coherent modes in the shallow water formalism, effectively silencing them within the inertial fluid.
@@ VI.3 Ferromagnetic magnons
In the inertial bulk, consisting of phase-locked agents that precess around the mean-field, the Goldstone mode yields spin waves in the form of ferromagnetic magnons. We derived the dispersion relation (Eq. 46),
$\omega(k)=\pm\frac{a_{0}}{R}k^{2}-i(a_{0}R)k^{2}.$
We make the following two key observations:
@@
The traditional hydrodynamic modes of transfer and the spintronic waves we propose to underpin information transfer are expected to coexist in a real chiral flock. Nevertheless, we expect both regimes in which orientation (magnons) should persist when waves in density (sound) fail, and regimes where physical transport (agent velocity) poses clear limitations on the second sound propagation, but where the magnons are nevertheless able to transmit phase gradients at great speed. These cases offer ways to test the degree to which a chiral active matter is indistinguishable from a dissipative spintronic fluid.
The above notwithstanding, in our model, propagating magnons cause a high-frequency smearing that rapidly delocalizes information. As shown in Figure 6, the dispersive nature of the magnon mode acts as an early-warning system, allowing high-frequency ‘panic’ signals to propagate through the marginally synchronized bulk well in advance of the density wavefront [39]. However, the marginal synchronization is likely realized on the spintronic fluid’s dissipative boundary, while in the ordered bulk $R\to 1$, the magnons are completely damped out. The spintronic chiral flock therefore does not turn or flow, but rather shatters and reforms, described in terms of shock dynamics in Ref. [8]. The Josephson inequality (Eq. 25) defines a threshold between trapped (superconducting, $|\omega-\dot{\Psi}|<I_{c}$) and phase-slipping (dissipative, $|\omega-\dot{\Psi}|>I_{c}$) turbulence. This shattering behaviour at impact is directly supported in the recent literature [40], and observations of “anomalous diffusion” during turns [39] are highly consistent with the notion of agents turning too abruptly, then phase-slip $|\omega-\dot{\Psi}|>I_{c}$. In the spintronic fluid, turns are initiated at the boundaries of the ordered flock where it costs almost no energy to phase-slip. This chirality is provided by the active bath ($\Delta\omega$), directly supported by Ref. [41]’s finding that turning flocks nucleate at the tips. What is more, the natural Goldstone dynamics with ferromagnetic magnons that we uncover in our model is consistent with observations of starling flocks where phase gradients travel faster than the bulk sound speed [41], as well as the paradigm of scale-free correlations in starling flocks [42].
Given the close correspondence between a spintronic fluid description and the various hydrodynamic descriptions that has accumulated, the tilted washboard potential should remain valid for many chiral active matter systems. It, and the Goldstone dynamics it drives, follow directly from the geometry of alignment itself, which, in the marginal synchronization regime, effectively forces the system to support such spin waves. Subsequently, aspects of chiral flocking, and the emergent Euler turbulence recovered in the model (Figure 1), are governed by the universality of the Landau-Lifshitz-Gilbert equation.
@@ VI.5 Topological quasiparticle gas
Ultimately, our discovery resolves the paradox of inertia in overdamped systems. In the hydrodynamic limit, Ref. [8] recovered inviscid Euler turbulence in the form of an Onsager dipole, necessitating a “phantom inertia.” In this article, we have shown that the propagation of Goldstone spin waves, caused by the azimuthal precession of the spintronic fluid, becomes mathematically indistinguishable from the advection of a massive fluid. That is, spin waves dragging phase texture along with them as they propagate constitute a topological quasi-particle.
@@ VI.6 Disorder as source of order
The key insight from our results lies in the origin of the description we have elucidated in this section. The 3D geometric constraints of polar, chiral particles lead directly to the two-dimensional motile Kuramoto model (Figure 4) and are entirely isomorphic to Gilbert damping in a magnetic field. The reason for this convergence is that both spintronics and chiral active matter obey the universality of the Landau-Lifshitz-Gilbert equation. Chiral active matter must, therefore, be strictly considered as a dissipative spintronic fluid, where the active bath at the boundaries provides the thermodynamic pressure that stabilizes the inertial bulk.

### v4

@@ Abstract
Recent minimalist modeling indicates that overdamped polar chiral active matter can support inviscid Euler turbulence, despite the system’s strictly dissipative microscopic nature. In this article, we establish the statistical mechanical foundation for this emergent inertial regime by deriving a formal isomorphism between the model’s agent dynamics and the overdamped Langevin equation for disordered Josephson junctions. We identify the trapped agent state as carrying non-dissipative phase rigidity supercurrents, a mapping we confirm empirically by demonstrating a disorder-broadened Adler-Ohmic crossover in the system’s slip velocity. Generalizing this framework to three dimensions ($S^{1}\to S^{2}$), we show that polar alignment on the unit sphere is geometrically equivalent to the Gilbert damping term in spintronics, and that the two-dimensional Kuramoto coupling term naturally appears in the tangent-plane projection of spin relaxation. This constraint generates inertial spin waves (ferromagnetic magnons) from the overdamped active bath, recovering the macroscopic transport predicted by Toner-Tu theory without invoking microscopic inertia. Our results indicate that chiral active matter should strictly be considered a dissipative spintronic fluid, where phase gradient transport is ensured by the Goldstone modes of the underlying broken symmetry.
@@ I Introduction
A recent minimalist, two-dimensional chiral active matter model offers a solution to the paradox by empirically demonstrating that an overdamped chiral flock can drive a conservative inertial cascade that culminates in inviscid Euler turbulence, in the form of a stable, sloshing Onsager dipole in the dynamic agent ensemble, where topological defects interact long-range through the Kirchoff-Onsager hamiltonian [8] (Figure 1). The present article provides detailed theoretical explanation for the findings in Ref. [8].
@@
The model in question is characterized by three simple rules, namely that its agents are (1) self-propelled, (2) polar, or flocking, favoring alignment, and (3) chiral, with intrinsic rotation [12, 8]. The model is unique in that it collapses these three properties into a single degree of freedom, phase, $\phi_{i}$,
@@
However, an important question remains. Our active agent ensemble is demonstrably overdamped and dissipative, yet the model exhibits an inverse, conservative energy cascade. How does a system that should fundamentally decay instead pump energy towards the largest scales? The resolution lies in considering the discrete distribution of agents as a spatially extended, disordered Josephson junction array. The depinning transition between ’running’ and ’trapped’ states then generates a macroscopic synchronization stiffness, or supercurrents of phase gradient information. By extending the model’s equations to three dimensions (3D), we eventually demonstrate that the rigidity supports Goldstone mode spin waves that map the two-dimensional (2D) model to the azimuthal precession around the mean-field of the 3D active chiral agents. We shall contrast these ferromagnetic magnons against the second sound waves of the Toner-Tu field theory, delineating the universality of the Landau-Lifshitz-Gilbert equation.
@@ IV Reconciling the model with real, three-dimensional chiral flocks
$\bm{\Omega}_{slip}=\underbrace{(\bm{\omega}_{i}-\bm{\Omega}_{\Psi})}_{\equiv\;\mathbf{I}_{\text{eff}}}-\underbrace{a_{0}R(\hat{\mathbf{\Psi}}\times\hat{\mathbf{n}}_{i})}_{\equiv\;\mathbf{I}_{c}}.$ (22)
@@
That is, the minimal agent dynamics mapping a two-dimensional unit phase vector to a three-dimensional unit vector naturally leads to an Adler relation. Remarkably, this occurs without the explicit definition of a Kuramoto-term. The reason for this stem from the recognition of the Landau-Lifshitz-Gilbert-equation from ferromagnetic theory, where a spin $\mathbf{S}$ naturally relaxes towards a magnetic field $\mathbf{H}$ (see Figure 4b). The three-dimensional synchronization in a chiral flock described in Eqs. (19) is therefore equivalent to magnetic damping.
@@ V Overdamped spin waves as ‘second sound’ in Toner-Tu theory
The previous section established a three-dimensional adherence to the Adler equations, where flock synchronization naturally arises from a Gilbert damping torque due by the geometry of the $S^{2}$ model (Eq. 19). Having established the rigidity of the flock, the natural next step is to demonstrate that the predicted effects are observable in real 3D flocks. To do so, we must reconcile the 3D extension of our minimalist model with the standard model invoked to explain flocking in three dimensions: the Toner-Tu field theory [9].
We begin with the 3D equations of motion, Eq. (19). To study the stability of the ordered state, we simplify by neglecting the intrinsic chirality $\bm{\omega}_{i}$, focusing on strong alignment:
@@
where $\xi$ is the characteristic interaction scale. This approximation follows from the definition of $\hat{\Psi}$ as the average director within an interaction radius $\sigma$. Treating the agent distribution as a continuum, $\hat{\Psi}(\mathbf{r})$ is the convolution of $\hat{n}(\mathbf{r})$ with a local weighting kernel. A Taylor expansion of $\hat{n}(\mathbf{r}+\bm{\delta})$ averaged over an isotropic neighborhood eliminates first-order gradient terms ($\langle\bm{\delta}\rangle=0$). The leading-order correction is therefore the quadratic term, proportional to $\nabla^{2}\hat{n}$, which encodes the local field curvature. Consequently, $\xi^{2}$ scales with the square of the interaction radius ($\xi\sim\sigma^{2}$).
Substituting the continuum relations into the alignment term of Eq. (26) yields:
@@
This form is physically significant as it mirrors the dissipative damping term $\hat{n}\times(\hat{n}\times\mathbf{H})$ found in the Landau-Lifshitz-Gilbert equation of spintronics [34].
To fully capture the wave-like transport observed in Toner-Tu theories [9], we restore the inertial term neglected in the overdamped limit. We introduce a coefficient of turning inertia, or precession stiffness, $\gamma$:
@@
We note that this restoration of effective inertia is physically mandated by the conservation of angular momentum density: as the local order parameter $R$ vanishes, the collective moment of inertia disappears ($I_{\text{eff}}\propto R$), compelling the precession rate to diverge to sustain a finite interaction torque. This description remains valid only while the local timescale separation persists, breaking down precisely at the defect cores and at the boundaries ($R\rightarrow 0$) where the divergence signals the transition from a hydrodynamic superfluid to a strictly microscopic, overdamped gas.
To satisfy dimensional consistency ($\gamma]=L^{2}T^{-1}$) within this physical constraint, we define the inertial coefficient as:
@@
which leads to a direct mapping to the sound speed of the topological gas, $c_{s}\approx v_{0}R$ [8],
@@
Physically, the inverse dependency in Eq. (31) implies that at the flock boundary ($R\to 0$), as alluded to, the precession frequency diverges. The system transitions from a “heavy,” locked superfluid in the bulk to a massless, fast-spinning gas. This divergence creates the active bath required to sustain the topological protection of the inner Onsager dipole [8].
@@
The active, 3D, overdamped bath on the boundaries of a real chiral flock naturally leads to a conservative azimuthal turning inertia of a spintronic fluid [34], by merit of the dimensional constraints of polar chiral alignment alone. The conservative fluid forms the characteristic flocking behaviour, and is strictly contingent on the agents individually adhering to the phase-locking condition from Josephson junction theory (Eq. 25). The conservation is the relaxation of the stress-energy of the sphere $S^{2}$, giving a geometric and thereby completely natural definition of synchronization in any system with polar alignment, by merit of the Landau-Lifshitz-Gilbert equation equation.
@@ VI.2 Thermodynamic pump
To see why the above holds, we identify the azimuthal angle of the $S^{2}$ potential as the scalar phase $\phi$ in the reduced $S^{1}$ limit. This degree of freedom is driven by the intrinsic chirality $\omega_{i}$ [and stochastic noise $\eta_{i}(t)$], providing the essential ”frustration” that prevents the chiral flock from freezing into a trivial, fully organized state [24, 8]. Instead, the frustration fuels a continuous, natural precession within the washboard potential wells (Figure 5b), generating the Goldstone mode quasiparticles (magnons) that sustain the fluid’s inertia.
@@
Inside the two-dimensional Onsager dipole’s vortex cores, we observe in Figure 2a,d) that the agents are continuously phase-slipping. In this limit, the critical current Eq. (12) vanishes, reducing the Adler equation to its trivial kinematic limit ($\dot{\phi}_{i}\approx\omega_{i}$), and the spintronic fluid’s capacity to support spin waves diverges, $\gamma\propto R^{-1}$ (Eq. 31). Consequently, the vortex cores are purely dissipative singularities where the capacity to sustain supercurrents has collapsed. This decoupling renders the cores unable to transmit phase stiffness or radiate coherent modes in the shallow water formalism, effectively silencing them within the inertial fluid [8].
@@ VI.3 Ferromagnetic magnons
In the 3D chiral flock, agents precess around the mean-field, and the Goldstone mode yields spin waves in the form of ferromagnetic magnons. We derived the dispersion relation (Eq. 46),
$\omega(k)=\pm\frac{a_{0}}{R}k^{2}-i(a_{0}R)k^{2},$
and we now make the following two key observations:
@@
The traditional hydrodynamic modes of transfer and the spintronic waves we propose to underpin information transfer are expected to coexist in a real chiral flock. Nevertheless, we expect regimes in which orientation (magnons) should persist when sound waves in density fail, and regimes where physical transport (agent velocity) poses clear limitations on the second sound propagation, but where the magnons are nevertheless able to transmit phase gradients at great speed. These cases offer ways to test the degree to which a chiral active matter is indistinguishable from a dissipative spintronic fluid.
The above notwithstanding, in our model, propagating magnons cause a high-frequency smearing that rapidly delocalizes information. As shown in Figure 6, the dispersive nature of the magnon mode acts as an early-warning system, allowing high-frequency ‘panic’ signals to propagate through the marginally synchronized bulk well in advance of the density wavefront [39]. However, the marginal synchronization is likely realized on the spintronic fluid’s dissipative boundary, while in the ordered bulk ($R\to 1$), the magnons are completely damped out. The spintronic chiral flock therefore does not turn or flow, but rather shatters and reforms, described in terms of shock merger in Ref. [8]. Here, the Josephson inequality (Eq. 25) defines a threshold between trapped (superconducting, $|\omega-\dot{\Psi}|<I_{c}$) and phase-slipping (dissipative, $|\omega-\dot{\Psi}|>I_{c}$) turbulence. This shattering behaviour at impact is directly supported in the literature [40], and observations of “anomalous diffusion” during turns [39] are highly consistent with the notion of agents turning too abruptly, then phase-slip $|\omega-\dot{\Psi}|>I_{c}$. In the spintronic fluid, turns are initiated at the boundaries of the ordered flock where it costs almost no energy to phase-slip. This chirality is provided by the active bath ($\Delta\omega$), directly supported by Ref. [41]’s finding that turning flocks nucleate at the tips. What is more, the natural Goldstone dynamics with ferromagnetic magnons that we uncover in our model is consistent with observations of starling flocks where phase gradients travel faster than the bulk sound speed [41], as well as the paradigm of scale-free correlations in starling flocks [42].
Given the close correspondence between a spintronic fluid description and the various hydrodynamic descriptions that has accumulated, the tilted washboard potential should remain valid for many chiral active matter systems. It, and the Goldstone dynamics it drives, follows directly from the geometry of alignment on the unit sphere, which, in the marginal synchronization regime, effectively forces the system to support such spin waves. Subsequently, aspects of chiral flocking, and the emergent Euler turbulence recovered in the model (Figure 1), are governed by the universality of the Landau-Lifshitz-Gilbert equation.
@@ VI.5 Topological quasiparticle gas
Ultimately, our discovery resolves the paradox of inertia in certain overdamped systems. In the hydrodynamic limit, Ref. [8] recovered inviscid Euler turbulence in the form of an Onsager dipole, necessitating a “phantom inertia.” In this article, we have shown that the propagation of Goldstone spin waves, caused by the azimuthal precession of the spintronic fluid, becomes mathematically indistinguishable from the advection of a massive fluid. That is, spin waves are dragging phase texture along as they propagate, constituting a topological quasi-particle.
@@ VI.6 Disorder as source of order
The key insight from our results lies in the origin of the description we have elucidated in this section. The 3D geometric constraints of polar, chiral agents lead directly to a two-dimensional Kuramoto coupling term in the tangent plane projection (Figure 4), and this action is entirely isomorphic to Gilbert damping in a magnetic field. The reason for this convergence is that both spintronics and chiral active matter obey the universality of the Landau-Lifshitz-Gilbert equation. Dry chiral active matter systens must, therefore, be strictly considered as dissipative spintronic fluids, where the active bath of the overdamped agents provides the thermodynamic pressure that stabilizes an inertial phase rigidity.